OpenStudy (anonymous):

Is x3 + x5 = x8? EXPLAIN your answer.

OpenStudy (unklerhaukus):

\[x^3 + x^5 = x^8\]?

OpenStudy (anonymous):

Mhm,

OpenStudy (lgbasallote):

Hint!!! \[\huge a^m \times a^n = a^{m+n}\] \[\huge a^m + a^n \ne a^{m+n}\]

OpenStudy (unklerhaukus):

\[x^3 + x^5 = x^8\] factorize out x^3 \[x^3(1 + x^2) = x^3(x^5)\] and cancel \[\cancel{x^3}(1 + x^2) = \cancel{x^3}(x^5)\]\[\qquad1 + x^2 = x^5\]

OpenStudy (unklerhaukus):

\[1+x^2=x^5\]take all terms to one side \[1+x^2-x^5=0\]

OpenStudy (anonymous):

Its eithers yes or no, Then explain why or why not

OpenStudy (unklerhaukus):

are there any solutions for x?

OpenStudy (anonymous):

ummmmm is the question above correct,yes or no?

OpenStudy (unklerhaukus):

is \[1+x^2−x^5=0\] true for x=1 for example?

OpenStudy (anonymous):

yeah thats easy AF your just combinding like terms

OpenStudy (anonymous):

so its correct @Suppa_Koopa_Troopa

OpenStudy (unklerhaukus):

as it turns out there are 5 solutions, four of them are complex, and one is real but irrational , in general \[x^3+x^5\neq x^8\]

OpenStudy (anonymous):

from the way i see it yeah unless its some parallel universe question lol .

OpenStudy (unklerhaukus):

\[1.194^3+1.194^5 \approx 1.194^8\]